(4 / x^6) ^3 I need to simplify it and I am pulling my hair out. Maybe I am too old for this! Can you possibly show me a detailed explanation?
Answer by stanbon(75887)

(4 / x^6) ^3 I need to simplify it
----------------------
= (4/x^6)(4/x^6)(4/x^6)
multiply the coefficients and add the exponents to get
= 64/x^18
================
Another way to do it:
(4/x^6)^3
= [4^3/x^(3*6)]
= 64/x^18
=====================
Cheers,
Stan H.